# Calogero–Sutherland hyperbolic system and Heckman–Opdam $\mathfrak{gl}_{n}$ hypergeometric function

### v1

@@ B.1 Proofs of lemmas
$\begin{split}\big{|}\check{I}_{r,J}\big{|}\leq C(g,\varepsilon)\,\hat{\Delta}({\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi\varepsilon\sum_{1\leq i\leq n-1-r}(x_{r+i}-x_{n})}.\end{split}$ (B.33)
@@
$\begin{split}\big{|}\check{\Phi}_{\bm{x}_{n}}(\bm{\lambda}_{n})-\check{\Phi}^{as}_{\bm{x}_{n}}(\bm{\lambda}_{n})\big{|}\leq C(g,\varepsilon)\,\hat{\Delta}({\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi\varepsilon\,m_{n}(\bm{x}_{n})},\end{split}$ (B.34)
@@
To do this, we use another combinatorial description of the linear order $\sigma$. Let us read the ordered set $\big{(}\{s_{1},\ldots,s_{m},t_{1},\ldots t_{m-1}\},\leavevmode\nobreak\ \leq\big{)}$ from left to right, starting from its minimal element. Meeting the first neighboring pair $t_{i_{1}}<s_{j_{1}}$ or $s_{j_{1}}<t_{i_{1}}$ we set
@@
$\begin{split}&\int_{Y_{i}}\prod_{k=1}^{K}(1+|t_{i}-s_{k}|)^{p_{k}}dt_{i}\leq\int_{s_{i}}^{s_{i+1}}\prod_{k=1}^{K}(1+|t_{i}-s_{k}|)^{p_{k}}dt_{i}\\[6.0pt] &\leq(s_{i+1}-s_{i})\,\prod_{s_{k}\leq s_{i}}(1+|s_{i+1}-s_{k}|)^{p_{k}}\,\prod_{s_{k}\geq s_{i+1}}(1+|s_{i}-s_{k}|)^{p_{k}}.\end{split}$ (B.44)
@@
$D\,\prod_{1\leq a<b\leq m}(1+|s_{a}-s_{b}|)^{A+2B+C}\,\prod_{k=1}^{n}(s_{i_{k+1}}-s_{i_{k}})\leq D\prod_{1\leq a<b\leq m}(1+|s_{a}-s_{b}|)^{A+2B+C+1},$ (B.52)
@@ Appendix C Evaluation at zero point
$\Psi_{\bm{\lambda}_{n}}(\bm{x}_{n})=\hat{\Lambda}_{n}(x_{n})\,\Psi_{\bm{\lambda}_{n-1}}(\bm{x}_{n-1})$ (C.53)
@@
$\Psi_{\lambda_{1}\lambda_{2}}(0)=\frac{1}{\Gamma(g)}\,\int_{\mathbb{R}}\prod_{k=1}^{2}\Gamma\bigl{(}\imath\lambda_{k}+{\textstyle\frac{g}{2}}-\imath\gamma\bigr{)}\Gamma\bigl{(}\imath\gamma-\imath\lambda_{k}+{\textstyle\frac{g}{2}}\bigr{)}\frac{d\gamma}{2\pi}=\frac{\Gamma(g)}{\Gamma(2g)}\,\Gamma(g+\imath\lambda_{12})\Gamma(g+\imath\lambda_{21})$
@@
$\Psi_{\lambda_{1}\lambda_{2}\lambda_{3}}(0)=\frac{\Gamma^{-2}(g)}{2!}\,\int_{\mathbb{R}^{2}}\,\frac{\prod_{i=1}^{3}\prod_{j=1}^{2}\,\Gamma\bigl{(}\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr{)}\,\Gamma\bigl{(}\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr{)}}{\Gamma(\imath\gamma_{12})\Gamma(\imath\gamma_{21})\Gamma(\imath\gamma_{12}+g)\Gamma(\imath\gamma_{21}+g)}\,\Psi_{\gamma_{1}\gamma_{2}}(0)\prod_{k=1}^{2}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\frac{1}{\Gamma(g)\Gamma(2g)}\,\frac{1}{2!}\,\int_{\mathbb{R}^{2}}\,\frac{\prod_{i=1}^{3}\prod_{j=1}^{2}\,\Gamma\bigl{(}\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr{)}\,\Gamma\bigl{(}\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr{)}}{\Gamma(\imath\gamma_{12})\Gamma(\imath\gamma_{21})}\,\prod_{k=1}^{2}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\frac{\Gamma(g)}{\Gamma(2g)}\,\frac{\Gamma(g)}{\Gamma(3g)}\,\prod_{1\leq i\not=j\leq 3}\,\Gamma(\imath\lambda_{ij}+g).$
@@
$\int_{\mathbb{R}^{n}}\,\frac{\prod_{k=1}^{n+1}\prod_{j=1}^{n}\Gamma(\alpha_{k}-\imath\lambda_{j})\Gamma(\imath\lambda_{j}-\beta_{k})}{\prod_{j<k}\Gamma(\imath\lambda_{k}-\imath\lambda_{j})\Gamma(\imath\lambda_{j}-\imath\lambda_{k})}\prod_{j=1}^{n}\frac{d\lambda_{j}}{2\pi}=\frac{n!\,\prod_{k,j=1}^{n+1}\Gamma(\alpha_{j}-\beta_{k})}{\Gamma\left(\sum_{j=1}^{n+1}(\alpha_{j}-\beta_{j})\right)}.$ (C.55)
@@
$\Psi_{\bm{\lambda}_{n+1}}(0)=\frac{\Gamma^{-n}(g)}{n!}\,\int_{\mathbb{R}^{n}}\,\frac{\prod_{i=1}^{n+1}\prod_{j=1}^{n}\,\Gamma\bigl{(}\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr{)}\,\Gamma\bigl{(}\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr{)}}{\prod_{1\leq i\not=j\leq n}\,\Gamma(\imath\gamma_{i}-\imath\gamma_{j})\Gamma(\imath\gamma_{j}-\imath\gamma_{i}+g)}\,\Psi_{\bm{\gamma}_{n}}(0)\prod_{k=1}^{n}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\Gamma^{-n}(g)\,\prod_{k=1}^{n}\frac{\Gamma(g)}{\Gamma(kg)}\,\frac{1}{n!}\,\int_{\mathbb{R}^{n}}\,\frac{\prod_{i=1}^{n+1}\prod_{j=1}^{n}\,\Gamma\bigl{(}\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr{)}\,\Gamma\bigl{(}\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr{)}}{\prod_{1\leq i\not=j\leq n}\,\Gamma(\imath\gamma_{i}-\imath\gamma_{j})}\,\prod_{k=1}^{n}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\Gamma^{-n}(g)\,\prod_{k=1}^{n}\frac{\Gamma(g)}{\Gamma(kg)}\,\frac{\Gamma^{n+1}(g)\,\prod_{1\leq i\not=j\leq n+1}\,\Gamma(\imath\lambda_{ij}+g)}{\Gamma((n+1)g)}=\prod_{k=1}^{n+1}\frac{\Gamma(g)}{\Gamma(kg)}\,\prod_{1\leq i\not=j\leq n+1}\,\Gamma(\imath\lambda_{ij}+g),$

### v3

@@ B.1 Proofs of lemmas
$\begin{split}\big|\check{I}_{r,J}\big|\leq C(g,\varepsilon)\,\hat{\Delta}({\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi\varepsilon\sum_{1\leq i\leq n-1-r}(x_{r+i}-x_{n})}.\end{split}$
@@
$\begin{split}\big|\check{\Phi}_{\bm{x}_{n}}(\bm{\lambda}_{n})-\check{\Phi}^{as}_{\bm{x}_{n}}(\bm{\lambda}_{n})\big|\leq C(g,\varepsilon)\,\hat{\Delta}({\bm{\lambda}}_{n})^{N(n)}\,e^{-2\pi\varepsilon\,m_{n}(\bm{x}_{n})},\end{split}$
@@
To do this, we use another combinatorial description of the linear order $\sigma$. Let us read the ordered set $\big(\{s_{1},\ldots,s_{m},t_{1},\ldots t_{m-1}\},\penalty 10000\ \leq\big)$ from left to right, starting from its minimal element. Meeting the first neighboring pair $t_{i_{1}}<s_{j_{1}}$ or $s_{j_{1}}<t_{i_{1}}$ we set
@@
$\begin{split}&\int_{Y_{i}}\prod_{k=1}^{K}(1+|t_{i}-s_{k}|)^{p_{k}}dt_{i}\leq\int_{s_{i}}^{s_{i+1}}\prod_{k=1}^{K}(1+|t_{i}-s_{k}|)^{p_{k}}dt_{i}\\[6.0pt] &\leq(s_{i+1}-s_{i})\,\prod_{s_{k}\leq s_{i}}(1+|s_{i+1}-s_{k}|)^{p_{k}}\,\prod_{s_{k}\geq s_{i+1}}(1+|s_{i}-s_{k}|)^{p_{k}}.\end{split}$
@@
$\begin{split}&D\cdot\prod_{1\leq a<b\leq m}(1+|s_{a}-s_{b}|)^{A+2B+C}\,\prod_{k=1}^{n}(s_{i_{k+1}}-s_{i_{k}})\leq\\ &D\cdot\prod_{1\leq a<b\leq m}(1+|s_{a}-s_{b}|)^{A+2B+C+1},\end{split}$ (B.37)
@@ Appendix C Evaluation at zero point
$\Psi_{\bm{\lambda}_{n}}(\bm{x}_{n})=\hat{\Lambda}_{n}(x_{n})\,\Psi_{\bm{\lambda}_{n-1}}(\bm{x}_{n-1})$
@@
$\Psi_{\lambda_{1}\lambda_{2}}(0)=\frac{1}{\Gamma(g)}\,\int_{\mathbb{R}}\prod_{k=1}^{2}\Gamma\bigl(\imath\lambda_{k}+{\textstyle\frac{g}{2}}-\imath\gamma\bigr)\Gamma\bigl(\imath\gamma-\imath\lambda_{k}+{\textstyle\frac{g}{2}}\bigr)\frac{d\gamma}{2\pi}=\frac{\Gamma(g)}{\Gamma(2g)}\,\Gamma(g+\imath\lambda_{12})\Gamma(g+\imath\lambda_{21}).$
@@
$\Psi_{\lambda_{1}\lambda_{2}\lambda_{3}}(0)=\frac{\Gamma^{-2}(g)}{2!}\,\int_{\mathbb{R}^{2}}\,\frac{\prod\limits_{i=1}^{3}\prod\limits_{j=1}^{2}\,\Gamma\bigl(\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr)\,\Gamma\bigl(\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr)}{\Gamma(\imath\gamma_{12})\Gamma(\imath\gamma_{21})\Gamma(\imath\gamma_{12}+g)\Gamma(\imath\gamma_{21}+g)}\,\Psi_{\gamma_{1}\gamma_{2}}(0)\prod\limits_{k=1}^{2}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\frac{1}{2\Gamma(g)\Gamma(2g)}\,\int\limits_{\mathbb{R}^{2}}\,\frac{\prod\limits_{i=1}^{3}\prod\limits_{j=1}^{2}\,\Gamma\bigl(\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr)\,\Gamma\bigl(\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr)}{\Gamma(\imath\gamma_{12})\Gamma(\imath\gamma_{21})}\,\prod_{k=1}^{2}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\frac{\Gamma(g)}{\Gamma(2g)}\,\frac{\Gamma(g)}{\Gamma(3g)}\,\prod_{1\leq i\not=j\leq 3}\,\Gamma(\imath\lambda_{ij}+g).$
@@
$\int_{\mathbb{R}^{n}}\,\frac{\prod\limits_{k=1}^{n+1}\prod\limits_{j=1}^{n}\Gamma(\alpha_{k}-\imath\lambda_{j})\Gamma(\imath\lambda_{j}-\beta_{k})}{\prod_{j<k}\Gamma(\imath\lambda_{k}-\imath\lambda_{j})\Gamma(\imath\lambda_{j}-\imath\lambda_{k})}\prod_{j=1}^{n}\frac{d\lambda_{j}}{2\pi}=\frac{n!\,\prod_{k,j=1}^{n+1}\Gamma(\alpha_{j}-\beta_{k})}{\Gamma\left(\sum_{j=1}^{n+1}(\alpha_{j}-\beta_{j})\right)}.$ (C.2)
@@
$\Psi_{\bm{\lambda}_{n+1}}(0)=\frac{\Gamma^{-n}(g)}{n!}\,\int\limits_{\mathbb{R}^{n}}\,\frac{\prod_{i=1}^{n+1}\prod_{j=1}^{n}\,\Gamma\bigl(\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr)\,\Gamma\bigl(\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr)}{\prod_{1\leq i\not=j\leq n}\,\Gamma(\imath\gamma_{i}-\imath\gamma_{j})\Gamma(\imath\gamma_{j}-\imath\gamma_{i}+g)}\,\Psi_{\bm{\gamma}_{n}}(0)\prod_{k=1}^{n}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\Gamma^{-n}(g)\,\prod_{k=1}^{n}\frac{\Gamma(g)}{\Gamma(kg)}\,\frac{1}{n!}\,\int_{\mathbb{R}^{n}}\,\frac{\prod\limits_{i=1}^{n+1}\prod\limits_{j=1}^{n}\,\Gamma\bigl(\imath\lambda_{i}-\imath\gamma_{j}+\frac{g}{2}\bigr)\,\Gamma\bigl(\imath\gamma_{j}-\imath\lambda_{i}+\frac{g}{2}\bigr)}{\prod_{1\leq i\not=j\leq n}\,\Gamma(\imath\gamma_{i}-\imath\gamma_{j})}\,\,\prod_{k=1}^{n}\frac{d\gamma_{k}}{2\pi}\\[6.0pt] =\Gamma^{-n}(g)\,\prod_{k=1}^{n}\frac{\Gamma(g)}{\Gamma(kg)}\,\frac{\Gamma^{n+1}(g)\,\prod_{1\leq i\not=j\leq n+1}\,\Gamma(\imath\lambda_{ij}+g)}{\Gamma((n+1)g)}=\prod_{k=1}^{n+1}\frac{\Gamma(g)}{\Gamma(kg)}\,\prod_{1\leq i\not=j\leq n+1}\,\Gamma(\imath\lambda_{ij}+g),$